One card is drawn from a well shuffled deck of 52 cards.

is equally likely, calculate the probability that the card will be

not an ace
Answer #1

We assume that the event 'Card drawn is an ace' is B

Therefore Card drawn is not an ace' should be B'.


We know that P(B) = 1- PB) =1-4/52=1-1/13=12/

13


answered by: Vijayan